OpenStudy (anonymous):

A chemist was asked to find the molecular weight of a gas and found that 3.73 g of the gas occupied 300 cm^3 at 25 celsius and 92.4 kpa. what is the molecular weight?

OpenStudy (anonymous):

ideal gas law

OpenStudy (anonymous):

you got the pressure as in kpa and the temperature in Centigrade and convert the volume of cm3 into m3.... and apply these three values in gas law equation...as PV = nRT.... n=PV/RT you can get the R value from internet.... then you will get n value... which is no. of moles.... so.. 1 mole = weight/ molecular weight.... so divide the given wight of sample with n value... you will get the mol. wt.

OpenStudy (somy):

connect 2 formulas Mole= mass/Mr PV= nRT n is mole so if we connect these 2 formulas we'll get PV= (mass/Mr)*RT u need to find Mr so now make Mr as a subject PV= (mass/Mr)*RT do cross multiplication and you'll get MrPV= mass*RT Mr= (mass*RT)/(PV) and now just substitute the values don't forget to convert cm3 to m3 (just divide by 10^6) , Celsius to Kelvin (to convert celsius to kelvin you just add celsius value to 273 and you'll get kelvin) , also change Kpa to pa (multiply by 1000) R is constant = 8.31 :)

OpenStudy (anonymous):

@Whitemonsterbunny17

OpenStudy (anonymous):

its ideal gas law sony kinda explain but I'm still confused

whitemonsterbunny17 (whitemonsterbunny17):

I'm sorry, I have no idea how to do this stuff, I wish I could've been more helpful :c

OpenStudy (anonymous):

thanks anyway

OpenStudy (anonymous):

@RANE

OpenStudy (anonymous):

sony tried to explain but she just confused me

OpenStudy (rane):

that needs time but currently im in class is it okay if i explain it later on?

OpenStudy (anonymous):

oh ok ill wait till ur done with class

OpenStudy (rane):

well, afetr this class i have another class so im basically finishing after almost 6hrs

OpenStudy (anonymous):

oh ok thanks anyway

OpenStudy (rane):

sorry

OpenStudy (anonymous):

@ParthKohli

OpenStudy (anonymous):

sony tried to explain but it just confused me

OpenStudy (somy):

@Karina1 do you know about formulas \[Mole = mass \div Mr\] \[PV = nRT\]